# AI Agents vs. Agentic AI: A Conceptual Taxonomy, Applications and Challenges

### v3

@@ VI Potential Solutions and Future Roadmap
The potential solutions (as illustrated in Figure 13) to these challenges and limitations of AI agents and Agentic AI are summarized in the following points:
Figure 13: Ten emerging architectural and algorithmic solutions such as RAG, tool use, memory, orchestration, and reflexive mechanisms addressing reliability, scalability, and explainability across both paradigms (AI Agents and Agentic AI)
@@
    Retrieval-Augmented Generation (RAG): For AI Agents, Retrieval-Augmented Generation mitigates hallucinations and expands static LLM knowledge by grounding outputs in real-time data [203]. By embedding user queries and retrieving semantically relevant documents from vector databases like FAISS Faiss or Pinecone Pinecone, agents can generate contextually valid responses rooted in external facts. This is particularly effective in domains such as enterprise search and customer support, where accuracy and up-to-date knowledge are essential.
    In Agentic AI systems, RAG serves as a shared grounding mechanism across agents. For example, a summarizer agent may rely on the retriever agent to access the latest scientific papers before generating a synthesis. Persistent, queryable memory allows distributed agents to operate on a unified semantic layer, mitigating inconsistencies due to divergent contextual views. When implemented across a multi-agent system, RAG helps maintain shared truth, enhances goal alignment, and reduces inter-agent misinformation propagation.
@@
    Tool-Augmented Reasoning (Function Calling): AI Agents benefit significantly from function calling, which extends their ability to interact with real-world systems [167, 204]. Agents can query APIs, run local scripts, or access structured databases, thus transforming LLMs from static predictors into interactive problem-solvers [162, 131]. This allows them to dynamically retrieve weather forecasts, schedule appointments, or execute Python-based calculations, all beyond the capabilities of pure language modeling.
    For Agentic AI, function calling supports agent level autonomy and role differentiation. Agents within a team may use APIs to invoke domain-specific actions such as querying clinical databases or generating visual charts based on assigned roles. Function calls become part of an orchestrated pipeline, enabling fluid delegation across agents [205]. This structured interaction reduces ambiguity in task handoff and fosters clearer behavioral boundaries, especially when integrated with validation protocols or observation mechanisms [14, 18].
@@
    Agentic Loop: Reasoning, Action, Observation: AI Agents often suffer from single-pass inference limitations. The ReAct pattern introduces an iterative loop where agents reason about tasks, act by calling tools or APIs, and then observe results before continuing. This feedback loop allows for more deliberate, context-sensitive behaviors. For example, an agent may verify retrieved data before drafting a summary, thereby reducing hallucination and logical errors. In Agentic AI, this pattern is critical for collaborative coherence. ReAct enables agents to evaluate dependencies dynamically reasoning over intermediate states, re-invoking tools if needed, and adjusting decisions as the environment evolves. This loop becomes more complex in multi-agent settings where each agent’s observation must be reconciled against others’ outputs. Shared memory and consistent logging are essential here, ensuring that the reflective capacity of the system is not fragmented across agents [132].
@@
    Memory Architectures (Episodic, Semantic, Vector): AI Agents face limitations in long-horizon planning and session continuity. Memory architectures address this by persisting information across tasks [206]. Episodic memory allows agents to recall prior actions and feedback, semantic memory encodes structured domain knowledge, and vector memory enables similarity-based retrieval [207]. These elements are key for personalization and adaptive decision-making in repeated interactions. Agentic AI systems require even more sophisticated memory models due to distributed state management. Each agent may maintain local memory while accessing shared global memory to facilitate coordination. For example, a planner agent might use vector-based memory to recall prior workflows, while a QA agent references semantic memory for fact verification. Synchronizing memory access and updates across agents enhances consistency, enables context-aware communication, and supports long-horizon system-level planning.
@@
    Multi-Agent Orchestration with Role Specialization: In AI Agents, task complexity is often handled via modular prompt templates or conditional logic. However, as task diversity increases, a single agent may become overloaded [208, 209]. Role specialization splitting tasks into subcomponents (e.g., planner, summarizer) allows lightweight orchestration even within single-agent systems by simulating compartmentalized reasoning. In Agentic AI, orchestration is central. A meta-agent or orchestrator distributes tasks among specialized agents, each with distinct capabilities. Systems like MetaGPT and ChatDev exemplify this: agents emulate roles such as CEO, engineer, or reviewer, and interact through structured messaging. This modular approach enhances interpretability, scalability, and fault isolation ensuring that failures in one agent do not cascade without containment mechanisms from the orchestrator.
@@
    Causal Modeling and Simulation-Based Planning: AI Agents often operate on statistical correlations rather than causal models, leading to poor generalization under distribution shifts. Embedding causal inference allows agents to distinguish between correlation and causation, simulate interventions, and plan more robustly. For instance, in supply chain scenarios, a causally aware agent can simulate the downstream impact of shipment delays. In Agentic AI, causal reasoning is vital for safe coordination and error recovery. Agents must anticipate how their actions impact others requiring causal graphs, simulation environments, or Bayesian inference layers. For example, a planning agent may simulate different strategies and communicate likely outcomes to others, fostering strategic alignment and avoiding unintended emergent behaviors. To enforce cooperative behavior, agents can be governed by a structured planning approach such as STRIPS or PDDL (Planning Domain Definition Language), where the environment is modeled with defined actions, preconditions, and effects. Inter-agent dependencies are encoded such that one agent’s action enables another’s, and a centralized or distributed planner ensures that all agents contribute to a shared goal. This unified framework supports strategic alignment, anticipatory planning, and minimizes unintended emergent behaviors in multi-agent systems.
@@
    Monitoring, Auditing, and Explainability Pipelines: AI Agents lack transparency, complicating debugging and trust. Logging systems that record prompts, tool calls, memory updates, and outputs enable post-hoc analysis and performance tuning. These records help developers trace faults, refine behavior, and ensure compliance with usage guidelines especially critical in enterprise or legal domains. For Agentic AI, logging and explainability are exponentially more important. With multiple agents interacting asynchronously, audit trails are essential for identifying which agent caused an error and under what conditions. Explainability pipelines that integrate across agents (e.g., timeline visualizations or dialogue replays) are key to ensuring safety, especially in regulatory or multi-stakeholder environments.
@@
Figure 14: Mindmap visualization of the future roadmap for AI Agents and Agentic AI.
@@
AI Agents are projected to evolve significantly through enhanced modular intelligence focused on five key domains as depicted in Figure 14 as : proactive reasoning, tool integration, causal inference, continual learning, and trust-centric operations. The first transformative milestone involves transitioning from reactive to Proactive Intelligence, where agents initiate tasks based on learned patterns, contextual cues, or latent goals rather than awaiting explicit prompts. This advancement depends heavily on robust Tool Integration, enabling agents to dynamically interact with external systems, such as databases, APIs, or simulation environments, to fulfill complex user tasks. Equally critical is the development of Causal Reasoning, which will allow agents to move beyond statistical correlation, supporting inference of cause-effect relationships essential for tasks involving diagnosis, planning, or prediction. To maintain relevance over time, agents must adopt frameworks for Continuous Learning, incorporating feedback loops and episodic memory to adapt their behavior across sessions and environments. Lastly, to build user confidence, agents must prioritize Trust & Safety mechanisms through verifiable output logging, bias detection, and ethical guardrails especially as their autonomy increases. Together, these pathways will redefine AI Agents from static tools into adaptive cognitive systems capable of autonomous yet controllable operation in dynamic digital environments.
Agentic AI, as a natural extension of these foundations, emphasizes collaborative intelligence through multi-agent coordination, contextual persistence, and domain-specific orchestration. Future systems (Figure 14 right side) will exhibit Multi-Agent Scaling, enabling specialized agents to work in parallel under distributed control for complex problem-solving mirroring team-based human workflows. This necessitates a layer of Unified Orchestration, where meta-agents or orchestrators dynamically assign roles, monitor task dependencies, and mediate conflicts among subordinate agents. Sustained performance over time depends on Persistent Memory architectures, which preserve semantic, episodic, and shared knowledge for agents to coordinate longitudinal tasks and retain state awareness. Simulation Planning is expected to become a core feature, allowing agent collectives to test hypothetical strategies, forecast consequences, and optimize outcomes before real-world execution. Moreover, Ethical Governance frameworks will be essential to ensure responsible deployment defining accountability, oversight, and value alignment across autonomous agent networks. Finally, tailored Domain-Specific Systems will emerge in fields like law, medicine, and supply chains, leveraging contextual specialization to outperform generic agents. This future positions Agentic AI not merely as a coordination layer on the top of AI Agents, but as a new paradigm for collective machine intelligence with adaptive planning, recursive reasoning, and collaborative cognition at its core.
@@
A transformative direction for future AI systems is introduced by the Absolute Zero: Reinforced Self-play Reasoning with Zero Data (AZR) framework, which reimagines the learning paradigm for AI Agents and Agentic AI by removing dependency on external datasets [217]. Traditionally, both AI Agents and Agentic AI architectures have relied on human-annotated data, static knowledge corpora, or preconfigured environments factors that constrain scalability and adaptability in open-world contexts. AZR addresses this limitation by enabling agents to autonomously generate, validate, and solve their own tasks, using verifiable feedback mechanisms (e.g., code execution) to ground learning. This self-evolving mechanism opens the door to truly autonomous reasoning agents capable of self-directed learning and adaptation in dynamic, data-scarce environments.
@@
In the context of Agentic AI—where multiple specialized agents collaborate within orchestrated workflows AZR lays the groundwork for agents to not only specialize but also co-evolve. For instance, scientific research pipelines could consist of agents that propose hypotheses, run simulations, validate findings, and revise strategies—entirely through self-play and verifiable reasoning, without continuous human oversight. By integrating the AZR paradigm, such systems can maintain persistent growth, knowledge refinement, and task flexibility across time. Ultimately, AZR highlights a future in which AI agents transition from static, pretrained tools to intelligent, self-improving ecosystems—positioning both AI Agents and Agentic AI at the forefront of next-generation artificial intelligence.
@@ VII Conclusion
In this study, we presented a comprehensive literature-based evaluation of the evolving landscape of AI Agents and Agentic AI, offering a structured taxonomy that highlights foundational concepts, architectural evolution, application domains, and key limitations. Beginning with a foundational understanding, we characterized AI Agents as modular, task-specific entities with constrained autonomy and reactivity. Their operational scope is grounded in the integration of LLMs and LIMs, which serve as core reasoning modules for perception, language understanding, and decision-making. We identified generative AI as a functional precursor, emphasizing its limitations in autonomy and goal persistence, and examined how LLMs drive the progression from passive generation to interactive task completion through tool augmentation.
@@
This study then explored the conceptual emergence of Agentic AI systems as a transformative evolution from isolated agents to orchestrated, multi-agent ecosystems. We analyzed key differentiators such as distributed cognition, persistent memory, and coordinated planning that distinguish Agentic AI from conventional agent models. This was followed by a detailed breakdown of architectural evolution, highlighting the transition from monolithic, rule-based frameworks to modular, role specialized networks facilitated by orchestration layers and reflective memory architectures. Additionally, this study then surveyed application domains in which these paradigms are deployed. For AI Agents, we illustrated their role in automating customer support, internal enterprise search, email prioritization, and scheduling. For Agentic AI, we demonstrated use cases in collaborative research, robotics, medical decision support, and adaptive workflow automation, supported by practical examples and industry-grade systems. Finally, this study provided a deep analysis of the challenges and limitations affecting both paradigms. For AI Agents, we discussed hallucinations, shallow reasoning, and planning constraints, while for Agentic AI, we addressed amplified causality issues, coordination bottlenecks, emergent behavior, and governance concerns. These insights offer a roadmap for future development and deployment of trustworthy, scalable agentic systems.
Acknowledgement

### v4

@@ VI Potential Solutions and Future Roadmap
VI-A Potential Solutions
To address the challenges and limitations of AI Agents and Agentic AI systems discussed in the previous section, we identify a set of promising solution pathways (as illustrated in Figure 13) including RAG , tool-augmented reasoning, memory architectures, causal modeling, reflexive mechanisms, orchestration frameworks, and governance-aware designs. These techniques collectively represent the frontier of efforts to overcome the brittleness, scalability bottlenecks, and coordination challenges that currently constrain both AI Agents and Agentic AI. At present, most deployed systems rely heavily on heuristic wrappers, manual prompt engineering, and shallow coordination logic, falling short of robust autonomy and reliability. In the following several paragraphs, we examine how each solution targets specific technical or systemic limitations, highlight gaps in current implementations, and propose future research directions to evolve these solutions into mature, interoperable components of next-generation intelligent systems. This roadmap is essential for transitioning from ad hoc agent deployments to principled, generalizable frameworks capable of powering scalable, safe, and context-aware agentic ecosystems.
Figure 13: Ten evolving architectural and algorithmic mechanisms such as RAG, tool augmentation, dynamic memory, causal modeling, orchestration, and reflexive self-evaluation are shown as key enablers to advance beyond prior usage toward addressing current limitations in reliability, scalability, and explainability. These techniques, while previously applied in isolated agent systems, are here recontextualized to support the demands of modern AI Agents and Agentic AI, enabling coordinated, adaptive, and verifiable behavior in increasingly complex and dynamic environments.
@@ VI-A Potential Solutions
    RAG : For AI Agents, RAG has the potential to mitigate hallucinations and can expand static LLM knowledge by grounding outputs in real-time data [219]. By embedding user queries and retrieving semantically relevant documents from vector databases like FAISS Source Link or Pinecone Pinecone, agents can generate contextually valid responses based on external facts. This retrieval-based grounding mechanism is particularly effective in domains such as enterprise search and customer support, where accuracy and access to up-to-date knowledge are essential for reliable task execution and user trust.
    In Agentic AI systems, RAG serves as a shared grounding mechanism across agents. For example, a summarizer agent may rely on the retriever agent to access the latest scientific papers before generating a synthesis. Persistent, queryable memory allows distributed agents to operate on a unified semantic layer, avoiding or minimizing inconsistencies due to divergent contextual views. When implemented across a multi-agent system, RAG helps maintain shared accuracy, enhances goal alignment, and reduces inter-agent misinformation propagation.
@@
    Tool-Augmented Reasoning (Function Calling): AI Agents benefit significantly from function calling, which extends their ability to interact with real-world systems [173, 220]. Agents can query APIs, run local scripts, or access structured databases, thus transforming LLMs from static predictors into interactive problem-solvers [168, 137]. This tool-augmented reasoning capability allows agents to dynamically access and process real-time, evolving information such as weather forecasts, stock prices, or user calendar updates and to perform executable actions like scheduling appointments, sending emails, or executing complex computations in Python. By bridging natural language reasoning with external tool interaction, this functionality empowers agents to go beyond static language generation and operate as autonomous, task-oriented decision-makers in real-world environments.
    For Agentic AI systems, function calling is instrumental in enhancing both autonomy and structured coordination among multiple agents. Each agent, assigned a specialized role within the system such as data retriever, visualizer, or decision-maker can independently invoke domain-specific APIs to perform targeted tasks, such as accessing clinical records or generating analytical dashboards. These function calls are not isolated; rather, they are embedded within an orchestrated pipeline a well-defined, multi-step workflow in which outputs from one agent seamlessly serve as inputs for the next [221]. This orchestration facilitates dynamic delegation, where agents can hand off subtasks based on predefined roles and capabilities without ambiguity or redundancy [21, 25].
    Moreover, integrating function calling within such orchestrated pipelines establishes clearer behavioral boundaries between agents. Each agent operates within its defined scope of responsibility, reducing the likelihood of overlapping actions or conflicting decisions. When coupled with validation protocols (e.g., response verification or schema checks in Waitgpt [222]) and observation mechanisms (e.g., feedback loops or audit logs [223, 224]), these boundaries are reinforced, ensuring that each agent not only performs its assigned task but does so transparently and accountably. This structured interaction model enhances system robustness, traceability, and ultimately the reliability of Agentic AI in complex, high-stakes domains.
@@
    Agentic Loop: Reasoning, Action, Observation: AI Agents often suffer from single-pass inference limitations [225]. The ReAct pattern introduces an iterative loop where agents reason about tasks, act by calling tools or APIs, and then observe results before continuing [138]. This feedback loop allows for more deliberate, context-sensitive behaviors. For example, an agent may verify retrieved data before drafting a summary, thereby reducing hallucination and logical errors. In Agentic AI, this pattern is critical for collaborative coherence. ReAct enables agents to evaluate dependencies dynamically reasoning over intermediate states, re-invoking tools if needed, and adjusting decisions as the environment evolves [138]. This loop becomes more complex in multi-agent settings where each agent’s observation must be reconciled against others’ outputs. Shared memory and consistent logging are essential here, ensuring that the reflective capacity of the system is not fragmented across agents.
@@
    Memory Architectures (Episodic, Semantic, Vector): As discussed before, AI Agents face limitations in long-horizon planning and session continuity. Memory architectures address this by persisting information across tasks [226]. Episodic memory allows agents to recall prior actions and feedback, semantic memory encodes structured domain knowledge, and vector memory enables similarity-based retrieval [227]. These elements are key for personalization and adaptive decision-making in repeated interactions. Agentic AI systems require even more sophisticated memory models due to distributed state management. Each agent may maintain local memory while accessing shared global memory to facilitate coordination. For example, a planner agent might use vector-based memory to recall prior workflows, while a QA agent references semantic memory for fact verification. Synchronizing memory access and updates across agents enhances consistency, enables context-aware communication, and supports long-horizon system-level planning.
@@
    Multi-Agent Orchestration with Role Specialization: In conventional AI Agent systems, increasing task complexity is often addressed through modular prompt engineering or conditional branching logic. However, as the range and intricacy of tasks grow, a single agent may become overburdened, leading to performance degradation or failure to generalize effectively [228, 229]. To mitigate this, role specialization dividing the overall task into discrete functional units such as planning, summarization, or verification enables a form of compartmentalized reasoning even within a single-agent framework. In this context, compartmentalized reasoning refers to the simulation of distinct cognitive functions within one agent by prompting it to reason through subtasks in sequence, often mimicking multiple expert roles.
    In contrast, Agentic AI systems institutionalize orchestration as a core architectural feature. Here, orchestration refers to the dynamic coordination and task delegation across a team of specialized agents, each designed to handle a specific sub-function in the overall workflow. This is typically governed by a meta-agent or orchestrator, a supervisory agent responsible for allocating tasks, managing dependencies, and maintaining global context across all agents. Systems like MetaGPT and ChatDev exemplify this paradigm: agents adopt predefined professional roles such as CEO, software engineer, or reviewer and communicate through structured messaging protocols to collaboratively complete complex projects. This orchestrated, role-specialized design enhances system interpretability by isolating reasoning traces within clearly defined agent roles. It also improves scalability, as tasks can be parallelized across agents, and contributes to fault tolerance, as errors from one agent are contained and monitored by the orchestrator, preventing systemic failure. Such modular, coordinated architectures are foundational to building robust and transparent Agentic AI systems.
@@
    Causal Modeling and Simulation-Based Planning: AI Agents often operate on statistical correlations rather than causal models, leading to poor generalization under distribution shifts. Embedding causal inference allows agents to distinguish between correlation and causation, simulate interventions, and plan counterfactually-informed, goal-directed actions that anticipate long-term effects and adapt to changing environments. For instance, in supply chain scenarios, a causally-aware agent can simulate the downstream impact of shipment delays. In Agentic AI, causal reasoning is vital for safe coordination and error recovery. Agents must anticipate how their actions impact others requiring causal graphs, simulation environments, or Bayesian inference layers. For example, a planning agent may simulate different strategies and communicate likely outcomes to others, fostering strategic alignment and avoiding unintended emergent behaviors. To enforce cooperative behavior, agents can be governed by a structured planning approach such as STRIPS or PDDL (Planning Domain Definition Language), where the environment is modeled with defined actions, preconditions, and effects. Inter-agent dependencies are encoded such that one agent’s action enables another’s, and a centralized or distributed planner ensures that all agents contribute to a shared goal. This unified framework supports strategic alignment, anticipatory planning, and minimizes unintended emergent behaviors in multi-agent systems.
@@
    Monitoring, Auditing, and Explainability Pipelines: AI Agents lack transparency, complicating debugging and trust. Logging systems that record prompts, tool calls, memory updates, and outputs enable post-hoc analysis and performance tuning. These records help developers trace faults, refine behavior, and ensure compliance with usage guidelines especially critical in enterprise or legal domains. Logging and explainability are even more critical for Agentic AI systems. With multiple agents interacting asynchronously, audit trails are essential for identifying which agent caused an error and under what conditions. Explainability pipelines that integrate across agents (e.g., timeline visualizations or dialogue replays) are key to ensuring safety, especially in regulatory or multi-stakeholder environments.
@@
Figure 14: Mindmap visualization of the future roadmap for AI Agents (left) and Agentic AI (right).
VI-B Future Roadmap
@@
AI Agents are projected to evolve significantly through enhanced modular intelligence focused on five key domains as depicted in Figure 14, which include proactive reasoning, tool integration, causal inference, continual learning, and trust-centric operations. The first transformative milestone involves transitioning from reactive to Proactive Intelligence, where agents initiate tasks based on learned patterns, contextual cues, or latent goals rather than awaiting explicit prompts. This advancement depends heavily on robust Tool Integration, enabling agents to dynamically interact with external systems, such as databases, APIs, or simulation environments, to fulfill complex user tasks. Equally critical is the development of Causal Reasoning, which will allow agents to move beyond statistical correlation, supporting inference of cause-and-effect relationships essential for tasks involving diagnosis, planning, or prediction. To maintain relevance over time, agents must adopt frameworks for Continuous Learning, incorporating feedback loops and episodic memory to adapt their behavior across sessions and environments. Lastly, to build user confidence, agents must prioritize Trust & Safety mechanisms through verifiable output logging, bias detection, and ethical guardrails especially as their autonomy increases. Together, these pathways will redefine AI Agents from static tools into adaptive cognitive systems capable of autonomous yet controllable operation in dynamic digital environments.
Agentic AI, as a natural extension of foundational AI agent frameworks, emphasizes collaborative intelligence through multi-agent coordination, contextual persistence, and domain-specific orchestration. Future systems (Figure 14, right side) are expected to exhibit Multi-Agent Scaling, enabling specialized agents to operate concurrently under distributed control to tackle complex, high-dimensional problems mirroring collaborative workflows typical of human teams. This scaling necessitates a layer of Unified Orchestration, wherein orchestrators specialized meta-agents assume responsibility for dynamically assigning roles, managing inter-agent communication, sequencing task dependencies, and resolving potential conflicts. Orchestration, in this context, refers to the intelligent coordination and regulation of interactions among multiple autonomous agents to ensure coherent and efficient collective behavior. Sustained performance over time will depend on robust Persistent Memory architectures that allow agents to store and retrieve semantic, episodic, and shared task-relevant knowledge, supporting continuity in longitudinal operations and enabling agents to maintain awareness of evolving goals and environmental states. Simulation Planning will become a core capability, empowering agent collectives to model hypothetical decision trajectories, forecast consequences, and optimize courses of action through internal trial-and-error mechanisms thus reducing real-world risk and increasing adaptive robustness.
Moreover, establishing Ethical Governance frameworks will be crucial to ensure that agent collectives operate within aligned moral and legal boundaries. These frameworks will define accountability structures, verification mechanisms, and safety constraints, especially in high-stakes domains involving autonomous decisions. Finally, the emergence of Domain-Specific Systems tailored for sectors such as law, medicine, logistics, and climate science will allow Agentic AI to leverage contextual specialization and outperform general-purpose agents through fine-tuned workflows and expert reasoning capabilities.
@@
A transformative direction for future AI systems is introduced by the Absolute Zero: Reinforced Self-play Reasoning with Zero Data (AZR) framework, which reimagines the learning paradigm for AI Agents and Agentic AI systems by removing dependency on external datasets [237]. Traditionally, both AI Agents and Agentic AI architectures have relied on human-annotated data, static knowledge corpora, or preconfigured environmental factors that constrain scalability and adaptability in open-world contexts. AZR addresses this limitation by enabling agents to autonomously generate, validate, and solve their own tasks, using verifiable feedback mechanisms (e.g., code execution) to ground learning. This self-evolving mechanism opens the door to truly autonomous reasoning agents capable of self-directed learning and adaptation in dynamic, data-scarce environments.
@@
In the context of Agentic AI where multiple specialized agents collaborate within orchestrated workflows, meaning structured, coordinated processes managed by a central controller or meta-agent AZR lays the groundwork for agents to not only specialize in distinct roles but also co-evolve through self-improving interactions and shared learning objectives [237]. For instance, scientific research pipelines could consist of agents that propose hypotheses, run simulations, validate findings, and revise strategies entirely through self-play and verifiable reasoning, without continuous human oversight. By integrating the AZR paradigm, such systems can maintain persistent growth, knowledge refinement, and task flexibility over time. Ultimately, AZR highlights a future in which AI agents transition from static, pretrained tools to intelligent, self-evolving and -improving ecosystems positioning both AI Agents and Agentic AI at the forefront of next-generation artificial intelligence.
@@ VII Conclusion
In this study, we presented a comprehensive literature-based evaluation of the evolving landscape of AI Agents and Agentic AI systems, offering a structured taxonomy that highlights foundational concepts, architectural evolution, application domains, and key limitations and potential solutions. Beginning with a foundational understanding, we characterized AI Agents as modular, task-specific entities with constrained autonomy and reactivity within the tasks specified for them. Their operational scope is enabled by the integration of LLMs and LIMs, which serve as core reasoning modules for perception, language understanding, and decision-making. We identified generative AI as a functional precursor to AI Agents, emphasizing its limitations in autonomy and goal persistence, and examined how LLMs drive the progression from passive generation to interactive task completion through tool augmentation.
This study then explored the conceptual emergence of Agentic AI systems as a transformative evolution from isolated agents or entities to orchestrated, multi-agent ecosystems that is, coordinated frameworks in which multiple specialized agents interact through structured role assignment, task delegation, and centralized or distributed control enabled by collaborative learning and collective decision making. We analyzed key differentiators such as distributed cognition, persistent memory, and coordinated planning that distinguish Agentic AI from conventional single-agent models. This analytical comparison was followed by a detailed breakdown of architectural evolution, highlighting the transition from monolithic, rule-based frameworks to modular, role-specialized networks. These networks are facilitated by orchestration layers, which serve as coordination mechanisms either centralized or decentralized that assign tasks, monitor agent interactions, and manage dependencies across specialized agents. Together with reflective memory architectures, these orchestration layers enable dynamic collaboration, task adaptability, and context preservation, marking a foundational shift toward scalable, intelligent agent collectives in Agentic AI systems.
@@
Additionally, this study surveyed application domains in which these two paradigms (AI Agents and Agentic AI systems) are deployed. For AI Agents, we illustrated their role in automating customer support, internal enterprise search, email prioritization, and scheduling. For Agentic AI, we showcased use cases in collaborative research, swarm robotics, medical decision support, and adaptive workflow automation, supported by practical examples and industry-grade systems. Finally, this study provided a deep analysis of the challenges and limitations affecting both paradigms. For AI Agents, we discussed hallucinations, shallow reasoning, and planning constraints as the key challenges faced, while for Agentic AI, we addressed amplified causality issues, coordination bottlenecks, emergent behavior, and governance concerns limiting the rapid advancement and adoption of these systems.
Through this comparative framework, we conclude that AI Agents serve well in narrow, tool-integrated scenarios with defined goals, while Agentic AI represents a paradigm shift toward distributed, multi-agent cognition capable of autonomous planning and decision-making. The transition from reactive task execution to orchestrated, collaborative workflows marks a significant milestone in the evolution of intelligent systems. These insights offer a roadmap for the future development and deployment of trustworthy, scalable Agentic AI systems that are capable of adapting to complex real-world environments
Acknowledgment